# The imprint of dark matter on the Galactic acceleration field

### v1

@@ Abstract
Measurements of the accelerations of stars enabled by time-series extreme-precision spectroscopic observations, from pulsar timing, and from eclipsing binary stars in the Solar Neighborhood offer insights into the mass distribution of the Milky Way that do not rely on traditional equilibrium modeling. Given the measured accelerations, we can determine a total mass density, and from this, by accounting for the mass in stars, gas, and dust, we can infer the amount of dark matter. Leveraging the FIRE-2 simulations of Milky Way-mass galaxies, we compare vertical acceleration profiles between cold dark matter (CDM) and self-interacting dark matter (SIDM) with constant cross-section of 1 cm2 g-1 across three halos with diverse assembly histories. Notably, significant asymmetries in vertical acceleration profiles near the midplane at fixed radii are observed in both CDM and SIDM, particularly in halos recently affected by mergers with satellites of Sagittarius/SMC-like masses or greater. These asymmetries offer a unique window into exploring the merger history of a galaxy. We show that SIDM halos consistently exhibit higher local stellar and dark matter densities and steeper vertical acceleration gradients, up to 30% steeper near the Solar Neighborhood. SIDM halos also manifest a more oblate halo shape in the Solar Neighborhood. Furthermore, enhanced precision in acceleration measurements and larger datasets promise to provide better constraints on the local dark matter density, complementing our understanding from kinematic analysis of their distribution within galaxies.
@@ 1 Introduction
This discrepancy has spurred exploration into alternative models, such as self-interacting dark matter (SIDM) (Carlson et al., 1992; Spergel & Steinhardt, 2000), which allows DM particles to interact with each other through elastic collisions. Such interactions can lead to a transfer of kinetic energy from the hot outer region to the dense inner region, resulting in a more cored density profile (Spergel & Steinhardt, 2000; Yoshida et al., 2000; Colín et al., 2002; Tulin & Yu, 2018). Moreover, a large self-interaction cross-section ( cm2 g-1) can induce a core-collapse phase that drives the DM mass towards the center and forms steeper inner DM profiles (Kahlhoefer et al., 2019; Zavala et al., 2019; Sameie et al., 2020).
@@
The most straightforward and least assumptive method for probing the mass distribution (including stars, gas and DM) in the Galaxy is through acceleration measurements of stars in the MW. Recent advances in high-precision spectrographs (e.g., Schwab et al., 2016; Fischer et al., 2016; Wright & Robertson, 2017; Pepe et al., 2021), improving precision in pulsar timing data (e.g., Keith et al., 2013; Pennucci, 2019; Goncharov et al., 2021), and high-precision spectroscopic observations of eclipsing binary stars by space-based missions (Hełminiak et al., 2019) now enable direct measurement of Galactic accelerations in the Solar Neighborhood with multiple independent techniques (e.g Silverwood & Easther, 2019; Chakrabarti et al., 2020, 2021; Phillips et al., 2021; Chakrabarti et al., 2022, 2022).
@@
These properties of the MW depend on the DM model (Sameie et al., 2018; Vargya et al., 2022; Sameie et al., 2021), indicating that measurement of the local dark matter density and the shape of the potential can provide discriminating power for constraining the nature of DM. Prior work on determinations of the Oort limit has focused on kinematic analysis (e.g. McKee et al., 2015; Schutz et al., 2018; Guo et al., 2020), which models a snapshot in time of the positions and speeds of stars under simplifying assumptions of equilibrium and symmetry both across the midplane and axisymmetrically, which can lead to inaccurate inferences of Galactic parameters for a time-dependent potential (Haines et al., 2019).
@@ 2.1 The Solar Neighborhood
Given the cosmological nature of these simulations in an arbitrary comoving box, we establish the galactocentric coordinates for each galaxy through a two-step process. First, we use the “shrinking spheres” method (Power et al., 2003) to determine the center position of each galaxy. We rotate the system to align the total angular momentum of the young stars (age Gyr) along the direction, which orients the Galactic disk in the XY plane for each simulation.111We also tested other methods to establish the disk plane, such as measuring the spatial midplane, and found that our results remain consistent regardless of the method used. We establish the Solar Circle with a fixed cylindrical radius of kpc (Gravity Collaboration et al., 2018). The rotation curves of these simulations are roughly flat at 8.1 kpc, with values close to km s-1, which roughly match that of the MW. Our choice is strictly motivated from the measured Solar position, in Sec. 3.2.1, we discuss our results for the DM density as a function of changing cylindrical radius from the center of each halo. While one could scale the cylindrical radius of each halo based on disk size or other parameters, this location would vary for each CDM and SIDM halo independently. Instead, we focus on what happens at a fixed radius away from the center in each halo.
We select a cylindrical region with a fixed Galactocentric cylindrical radius of kpc and a height of kpc, centered around the Solar Circle. This cylindrical region is then divided into small 50 kpc bins222This value is times larger than the star and gas softening parameters used in the simulations and about 2 times greater than the DM softening. in the X and Y direction, covering the kpc. Within each of these bins, we independently compute the surface density () for baryons (stars and gas), DM, and for all (baryons and DM) the species combined together. Following this, we introduce a metric, , to quantify the relative surface density variation:
@@
Next, we categorize the bins into three density regions – low, median, and high – based on their surface density in quartile ranges percentile, percentile, and percentile respectively. The three density regimes effectively provide upper and lower limits on the Galactic acceleration profiles.
@@
Table LABEL:tab:gal_prop summarizes the galaxy-wide properties along with average variations in the low, median, and high density regions around the Solar Circle for all of the simulations.333Varying ranges to match the MW estimates. While, the median baryonic volume density in the Solar Circle across all halos is generally much lower () compared to the MW, the total surface density across the halos is relatively close to that of the MW (within ), except for m12i CDM that is about away. Also, McCluskey et al. (2024) motivated that the MW forms an unusually dynamically cold disk which could explain higher average density, contrasting with other observed galaxies of similar mass ( M) through analysis of the Latte disks. Additionally, our estimates for the MW are solely based on the local volume around the Solar Neighborhood. In terms of matching local properties, the m12m SIDM halo is the most similar to the MW, exhibiting comparable thin and thick disk, and cold gas scale heights.
We observe that both the total volume and surface density between regions within a same halo can vary by a factor 2. For example, in the low, median, and high density regions, the volume density444for kpc to match the volume density calculation in the MW for m12m SIDM is approximately (36, 50, 79) . Only of our selected bins exhibit total volume densities within of the MW value of and 10% bins have total volume density values higher than the MW. The majority of the variations arise from the baryonic component (stars and gas) and the DM component is evenly distributed azimuthally (lower errors in Table LABEL:tab:gal_prop, see Fig. 1.).
@@
Comparing DM models, the SIDM halos exhibit both higher stellar and DM volume and surface densities in the Solar Neighborhood than their CDM counterparts. This is primarily due to SIDM particles’ capacity to exchange energy and momentum in addition to gravity, making them more responsive and sensitive to the baryonic potential (Sameie et al., 2018; Elbert et al., 2018; Despali et al., 2019; Robles et al., 2019; Santos-Santos et al., 2020). The higher DM density facilitates the entrapment of additional gas, forming new stars and establishing a feedback loop, where DM particles can respond faster to the rising stellar density (Despali et al., 2019; Sameie et al., 2021). While at about 8.1 kpc from the Galactic center, we anticipate only about one DM scattering event per Hubble time, these rates are higher in the inner regions of the galaxy (Vargya et al., 2022), which leads to more pronounced differences in DM density between CDM and SIDM in the inner regions (see Fig. 5 in Sec. 3.2.1). Consequently, SIDM halos exhibit higher density in DM and stellar distribution within the baryon-dominated potential, resulting in increased oblateness in the inner regions and around the Solar Circle.
@@ 2.2 Comparison with local potential models
Ideally, one would integrate the acceleration field (from e.g., pulsar timing or other acceleration measurements) to infer the potential directly. However, in practice, the scarcity of data points (e.g., 20 pulsars within 2 kpc of the Sun (Donlon et al., 2024)) currently limits our ability to perform such integration. Consequently, in the MW, observes use static parameterize models to fit the available data accurately. Here, we present our simulations with similar parameters through potential model fitting on stars in order for a direct comparison. We consider two axisymmetric potentials previously used in Chakrabarti et al. (2021). Specifically, a low-order (LO) expansion potential (eq. 2, Chakrabarti et al., 2021) near the position of the sun (i.e adopted Solar Circle in our case), and a single Miyamoto-Nagai (MN) potential (eq. 3) for the Galactic disk.
@@
where is the local standard of rest velocity, determines the shape of the potential affecting the vertical accelerations, and represents the squared angular frequency of oscillations in the vertical density profile. A higher value of value indicates a flatter potential profile in the Solar Neighborhood. F
@@
Table 2 shows the best fit parameters and oblateness () obtained for the simulations using the LO and MN potentials along with the best fit parameters for the MW from Donlon et al. (2024). It is noteworthy that the values derived from the LO potential () and the MN disk () are similar in all simulations, except for the m12m SIDM case. This suggests that the shape of the local potential at the Solar Circle is primary influenced by the stellar disk.
@@ 3 Vertical acceleration profiles
We calculate the vertical accelerations above and below the Galactic midplane using a binned method ( kpc) to present smooth trends as a function of vertical distance () from the midplane.
We slice the low, median, and high density regions as defined in Sec. 2.1 into 50 kpc vertical distance bins555This value is times greater than the softening parameters used in the simulations.. Next, we average the gravitational accelerations computed for stars and DM particles in each region within each bin. With this approach, we obtain a smooth and accurate representation of the median acceleration profile; the high density bin provide an upper limit, and the low density bin provide a lower limit for each profile.
Fig. 2 plots the vertical acceleration (top row) and change in the vertical velocity over a 10-year baseline () (bottom row) at present day for the m12i (left column), m12f (middle column), and m12m (right column) in CDM (blue) and SIDM (green) model. The shaded regions represent the upper and lower limits on the profiles derived from the high and low density regions respectively. SIDM halos have consistently larger magnitude of their accelerations than CDM, thanks to their higher surface densities, especially at larger distances from the mid-plane. We note that the degree of difference in between SIDM and CDM profiles varies with the simulation, with the smallest difference observed for m12f (one with a massive merger about 3 Gyr ago before the present day) and the largest difference for m12m (no mergers in the last 10 Gyr). This observation suggests that mergers may have a role in removing or mitigating differences between SIDM and CDM profiles. We discuss this more in Sec. 3.2.
@@ 4 Discussion and conclusions
The shape of the potential and density in the Solar Neighborhood are predominantly influenced by the Galactic disk (see Sec. 2.2). However, SIDM particles greater responsiveness to the disk potential leads to higher flattening in the DM shape at Solar Circle. This leads to measurable distinctions in the local dark matter density within the Solar Neighborhood. SIDM halos consistently demonstrate denser and more oblate potentials, resulting in vertical acceleration gradient profiles that are steeper by 10-30% compared to CDM. (see Sec. 3.1 and Fig. 3). However, the galaxy-to-galaxy variation in density is broader than the difference between SIDM and CDM at 1 cm2/g.

### v2

@@ Abstract
Measurements of the accelerations of stars enabled by time-series extreme-precision spectroscopic observations, pulsar timing, and eclipsing binary stars in the Solar Neighborhood offer insights into the mass distribution of the Milky Way that do not rely on traditional equilibrium modeling. Given the measured accelerations, we can determine a total mass density and infer the amount of dark matter (DM) by accounting for the mass in stars, gas, and dust. Leveraging FIRE-2 simulations of Milky Way-mass galaxies, we compare vertical acceleration profiles between cold DM (CDM) and self-interacting DM (SIDM) with constant cross-section of 1 cm2g-1 across three halos with diverse assembly histories. Notably, significant asymmetries in vertical acceleration profiles near the midplane at fixed radii are observed in both CDM and SIDM, particularly in halos recently affected by mergers with satellites of Sagittarius/SMC-like masses or greater. These asymmetries offer a unique window into exploring the merger history of a galaxy. We show that SIDM halos manifest a more oblate shape and consistently exhibit higher local stellar and DM densities and steeper vertical acceleration gradients, up to 10-30% steeper near the Solar Neighborhood. However, similar magnitude changes can arise from azimuthal variations in the baryonic components at a fixed radius and external influences like mergers, making it difficult to distinguish between CDM and SIDM using acceleration measurements in a single galaxy.
@@ 1 Introduction
This discrepancy has spurred exploration into alternative models, such as self-interacting dark matter (SIDM) (Carlson et al., 1992; Spergel & Steinhardt, 2000), which allows DM particles to interact with each other through elastic collisions. Such interactions can lead to a transfer of kinetic energy from the hot outer region to the dense inner region, resulting in a more cored density profile (Spergel & Steinhardt, 2000; Yoshida et al., 2000; Colín et al., 2002; Gentile et al., 2004; Tulin & Yu, 2018; Salucci, 2019). Moreover, a large self-interaction cross-section ( cm2 g-1) can induce a core-collapse phase that drives the DM mass towards the center and forms steeper inner DM profiles (Kahlhoefer et al., 2019; Zavala et al., 2019; Sameie et al., 2020).
@@
The most straightforward and least assumptive method for probing the mass distribution (including stars, gas and DM) in the Galaxy is through acceleration measurements of stars in the MW. Recent advances in high-precision spectrographs (e.g., Schwab et al., 2016; Fischer et al., 2016; Wright & Robertson, 2017; Pepe et al., 2021), improving precision in pulsar timing data (e.g., Keith et al., 2013; Pennucci, 2019; Goncharov et al., 2021), and high-precision spectroscopic observations of eclipsing binary stars by space-based missions (Hełminiak et al., 2019) now enable direct measurement of Galactic accelerations in the Solar Neighborhood with multiple independent techniques (e.g Quercellini et al., 2008; Silverwood & Easther, 2019; Chakrabarti et al., 2020, 2021; Phillips et al., 2021; Chakrabarti et al., 2022, 2022).
@@
These properties of the MW depend on the DM model (Sameie et al., 2018; Vargya et al., 2022; Sameie et al., 2021), indicating that measurement of the local dark matter density and the shape of the potential can provide discriminating power for constraining the nature of DM. Prior work on determinations of the Oort limit has focused on kinematic analysis (e.g. Salucci et al., 2010; McKee et al., 2015; Schutz et al., 2018; Guo et al., 2020; de Salas & Widmark, 2021), which models a snapshot in time of the positions and speeds of stars under simplifying assumptions of equilibrium and/or symmetry both across the midplane and axisymmetrically, which can lead to inaccurate inferences of Galactic parameters for a time-dependent potential (Haines et al., 2019).
@@ 2.1 The Solar Neighborhood
Given the cosmological nature of these simulations in an arbitrary comoving box, we establish the galactocentric coordinates for each galaxy through a two-step process. First, we use the “shrinking spheres” method (Power et al., 2003) to determine the center position of each galaxy. We rotate the system to align the total angular momentum of the young stars (age Gyr) along the direction, which orients the Galactic disk in the XY plane for each simulation.111We also tested other methods to establish the disk plane, such as measuring the spatial midplane, and found that our results remain consistent regardless of the method used.
We establish the Solar Circle with a fixed cylindrical radius of kpc (Gravity Collaboration et al., 2018). The rotation curves of these simulations are roughly flat at 8.1 kpc, with values close to km s-1, which roughly match that of the MW. Our choice is strictly motivated from the measured Solar position, in Sec. 3.2.1, we discuss our results for the DM density as a function of changing cylindrical radius from the center of each halo and show that the DM densities at kpc are consistent with the MW values, within 1 (see Table LABEL:tab:gal_prop). Although one could scale the cylindrical radius of each halo based on quantities like the exponential scale length of the Galactic disk or the stellar half-mass radius, the galaxies in this paper are within a factor of 2 in stellar mass (see Table LABEL:tab:gal_prop). Thus, scaling to a specific scale radius is less critical compared to cases where galaxies have widely different masses. The decision to scale depends on the properties being probed and whether there is a valid reason to expect them to scale with the scale radius at a fixed mass. For example, Bellardini et al. (2021) found that measuring FIRE-2 galaxies in physical units rather than scaling to a specific radius resulted in less scatter and more self-similarity in their analysis of metallicity radial gradients. Since, our results depend on the density at a fixed radius, which does not necessarily scale with the scale radius at a fixed mass, the meaningful “scaling” in this context is to match the local densities rather than adjusting based on a scale radius. We find that using a fixed radius from the center of each halo offers a more consistent and reliable comparison for our analysis.
We select a cylindrical region with a fixed Galactocentric cylindrical radius of kpc and a height of kpc, centered around the Solar Circle. This cylindrical region is then divided into small bins222This value is times larger than the star and gas softening parameters used in the simulations and about 2 times greater than the DM softening. of 50 kpc in the X and Y direction, covering the kpc. Within each of these bins, we independently compute the surface density () for baryons (stars and gas), DM, and for all (baryons and DM) the species combined together. Following this, we introduce a metric, , to quantify the relative surface density variation:
@@
Next, we categorize the bins into three density regions – low, median, and high – based on their surface density in quartile ranges percentile, percentile, and percentile respectively. The three density regimes effectively provide upper and lower limits on the Galactic acceleration profiles. Table LABEL:tab:gal_prop summarizes the galaxy-wide properties along with average variations in the low, median, and high density regions around the Solar Circle for all of the simulations.333Varying ranges to match the MW estimates. While, the median baryonic volume density in the Solar Circle across all halos is generally much lower () compared to the MW, the total surface density across the halos is relatively close to that of the MW (within ), except for m12i CDM that is about away. Also, McCluskey et al. (2024) showed that the MW forms an unusually dynamically cold disk which could explain higher average density, contrasting with other observed galaxies of similar mass ( M) through analysis of the Latte disks. Additionally, our estimates for the MW are solely based on the local volume around the Solar Neighborhood. In terms of matching local properties, the m12m SIDM halo is the most similar to the MW, exhibiting comparable thin and thick disk, and cold gas scale heights. We observe that both the total volume and surface density between regions within a same halo can vary by a factor 2. For example, in the low, median, and high density regions, the volume density444for kpc to match the volume density calculation in the MW for m12m SIDM is approximately (36, 50, 79) . Only of our selected bins exhibit total volume densities within of the MW value of and 10% bins have total volume density values higher than the MW. The majority of the variations arise from the baryonic component (stars and gas) and the DM component is evenly distributed azimuthally (lower errors in Table LABEL:tab:gal_prop, see Fig. 1.).
@@
Comparing DM models, the SIDM halos exhibit both higher stellar and DM volume and surface densities in the Solar Neighborhood than their CDM counterparts. This is primarily due to SIDM particles’ capacity to exchange energy and momentum in addition to gravity, making them more responsive and sensitive to the baryonic potential (Sameie et al., 2018; Elbert et al., 2018; Despali et al., 2019; Robles et al., 2019; Santos-Santos et al., 2020). The higher DM density facilitates the entrapment of additional gas, forming new stars and establishing a feedback loop, where DM particles can respond faster to the rising stellar density (Despali et al., 2019; Sameie et al., 2021). While at about 8.1 kpc from the Galactic center, we anticipate only about one DM scattering event per particle per Hubble time, these rates are higher in the inner regions of the galaxy (Vargya et al., 2022), which leads to more pronounced differences in DM density between CDM and SIDM in the inner regions (see Fig. 5 in Sec. 3.2.1). Consequently, SIDM halos exhibit higher density in DM and stellar distribution within the baryon-dominated potential, resulting in increased oblateness in the inner regions and around the Solar Circle.
@@ 2.2 Comparison with local potential models
Ideally, one would integrate the acceleration field (from e.g., pulsar timing or other acceleration measurements) to infer the potential directly. However, in practice, the scarcity of data points (e.g., 20 pulsars within 2 kpc of the Sun (Donlon et al., 2024)) currently limits our ability to perform such integration. Consequently, in the MW, observers use static parameterize models to fit the available data accurately. Here, we present our simulations with similar parameters through potential model fitting on stars in order for a direct comparison. We consider two axisymmetric potentials previously used in Chakrabarti et al. (2021). Specifically, a low-order (LO) expansion potential (eq. 2, Chakrabarti et al., 2021) near the position of the sun (i.e adopted Solar Circle in our case), and a single Miyamoto-Nagai (MN) potential (eq. 3) for the Galactic disk.
@@
where is the local standard of rest velocity, determines the shape of the potential affecting the vertical accelerations, and represents the squared angular frequency of oscillations in the vertical density profile. A higher value of value indicates a flatter potential profile in the Solar Neighborhood.
@@
Table 2 shows the best fit parameters and oblateness () obtained for the simulations using the LO and MN potentials along with the best fit parameters for the MW from Donlon et al. (2024). The scale radius (), and scale height () from the MN profile fits for all three halos are within 1 of each other when comparing the CDM and SIDM models. It is noteworthy that the values derived from the LO potential () and the MN disk () are similar in all simulations, except for the m12m SIDM case. This suggests that the shape of the local potential at the Solar Circle is primary influenced by the stellar disk.
@@ 3 Vertical acceleration profiles
We calculate the vertical accelerations above and below the Galactic midplane using a binned method ( kpc) to present smooth trends as a function of vertical distance () from the midplane. We slice the low, median, and high density regions as defined in Sec. 2.1 into 50 kpc vertical distance bins555This value is times greater than the softening parameters used in the simulations.. Next, we average the gravitational accelerations computed for stars and DM particles in each region within each bin. With this approach, we obtain a smooth and accurate representation of the median acceleration profile; the high density bin provide an upper limit, and the low density bin provide a lower limit for each profile.
Fig. 2 plots the vertical acceleration (top row) and change in the vertical velocity over a 10-year baseline () (bottom row) at present day for the m12i (left column), m12f (middle column), and m12m (right column) in CDM (blue) and SIDM (red) model. The shaded regions represent the upper and lower limits on the profiles derived from the high and low density regions respectively. SIDM halos have consistently larger magnitude of their accelerations than CDM, thanks to their higher surface densities, especially at larger distances from the mid-plane. The error bars in the first column represent current average measurement uncertainties for pulsar accelerations up to kpc (gray error bars) centered on the m12i SIDM value at kpc for reference, derived from pulsar timing data over a 10-year baseline, and the expected uncertainties for a 20-year baseline (green error bars) as reported in Donlon et al. (2024) scaling pulsar timing precision by a power of 5/3 (Lorimer & Kramer, 2005). The estimated future uncertainties are expected to distinguish between the profile differences noted between the CDM and SIDM models. We note that the degree of difference in between SIDM and CDM profiles varies with the simulation, with the smallest difference observed for m12f (one with a massive merger about 3 Gyr ago before the present day) and the largest difference for m12m (no mergers in the last 10 Gyr). This observation suggests that mergers may have a role in removing or mitigating differences between SIDM and CDM profiles. We discuss this more in Sec. 3.2.
@@ 4 Discussion and conclusions
The shape of the potential and density in the Solar Neighborhood are predominantly influenced by the Galactic disk (see Sec. 2.2). However, SIDM particles show greater responsiveness to the disk potential leads to higher flattening in the DM shape at Solar Circle. This leads to measurable distinctions in the local dark matter density within the Solar Neighborhood. SIDM halos consistently demonstrate denser and more oblate potentials, resulting in vertical acceleration gradient profiles that are steeper by 10-30% compared to CDM. (see Sec. 3.1 and Fig. 3). However, the galaxy-to-galaxy variation in density is broader than the difference between SIDM and CDM at 1 cm2/g.